OpenStudy (anonymous):

Appleton’s population is 400 greater than Berryville’s population. If Berryville’s population were reduced by 900 people, then Appleton’s population would be 3 times as large as Berryville’s population. What is Berryville’s current population?

imqwerty (imqwerty):

let the current population of berryville be x and that of appleton be y so y=400+x 1st equation y=3(x-300) 2nd equation hope this helps :)

imqwerty (imqwerty):

nw u jst have to solve equation 1 and 2 to get the values of x and y.

OpenStudy (anonymous):

but how is 400+x equal to 3(x-300)?

OpenStudy (anonymous):

it says it's 400 greater not 3 times greater

imqwerty (imqwerty):

400+x = y and y=3(x-300) these two r different things i'll elaborate according to the statement -Appleton’s population is 400 greater than Berryville’s population---- we took Appleton's population as y and berryville's popu. as x and made the equation :- y=400+x according to the statement- If Berryville’s population were reduced by 900 people, then Appleton’s population would be 3 times as large as Berryville’s population-------we made the second equation :- y=3(x-300) nw u hav to equate these two equations to get the answer

OpenStudy (anonymous):

yea that's what im saying don't they have to equal each other in order to make it 400+x=3(x-300)

OpenStudy (anonymous):

Did every one else get 1550 for Berryville’s population?

imqwerty (imqwerty):

yes u r correct